# Extracting Affect Aggregates from Longitudinal Social Media Data with Temporal Adapters for Large Language Models

### v1

@@ Results
We implement two additional survey questions besides the one used in Britain’s Mood, Measured Weekly. First, we focus on demonstrating the robustness of our extraction method and extract the collective emotions scared and sad the PANAS-X survey instrument. The instrument works quite differently from YouGov’s questionnaire in that it measures each emotion with multiple adjectives, with the same answer options for each adjective: ‘very slightly or not at all’, ‘a little’, etc. Again, we cross-correlate the extracted collective emotions with the YouGov survey data – Figure 6(a) shows the results from 3 training seeds across the different extraction methods. We find that both scared and sad achieve cross-correlations comparable to the ones we extracted with the original YouGov question wording. This supports the robustness of our extraction method across prompt formulations, i.e., survey instruments, when measuring the same construct.
@@
Second, we go beyond measuring collective emotions and tested a question that measures attitudes towards the National Health Service (NHS). YouGov started running their weekly and monthly surveys on different days in May 2020, so we train 3 additional adapters to extract attitudes in exactly the same weeks that the monthly survey was run. YouGov runs most of its ‘trackers’ only in monthly waves, including on the attitude towards the NHS. This means that during the time period that we study, we only have 9 data points available. Here we see a potential clear benefit of extracting attitudes from social media data – it is much less expensive and can have a higher temporal resolution. Figure 6(b) shows the cross-correlation of the two available answer options for the 9 available data points between our estimates and the survey data – for time series, see Figure 10 in the Appendix. While we find positive correlations, they are not significant across all random seeds and the different seeds create a large confidence interval. Reasons for this might be measurement error in our extraction method for a more complex construct, and less discussion around this particular topic in our training data. Still, this result shows that our method can, in principle, be applied to the extraction of more complex attitudes as well.
Discussion
@@
Our method focuses on the extraction of separate time series data, rather than comparing the prevalence of moods or attitudes cross-sectionally. The reason is that the relative probabilities of answer options in a cross-sectional sense are strongly influenced by LLM pre-training. Future research should combine these, so far orthogonal, approaches by correcting for differences that stem from model pre-training. The results we present were obtained by training weekly Temporal Adapters, following exactly the intervals in which the respective survey waves were conducted. This allowed for improved comparability between our estimates and survey data, but higher temporal resolution, a potential benefit of using social media data, should be investigated in future research. In regions where survey data is available, survey data could in the future be used as a prior for predicting affect aggregates with higher temporal resolution from social media data. Yet, our method remains applicable to regions in which no (recent) survey data is available, e.g., because of lack of infrastructure or an ongoing violent conflict. Like any other attitude extraction method, our approach requires validation in additional contexts, with possible design decisions ranging from model training hyperparameters to the exact question wording and the answer extraction method. Future research should further investigate the applicability of this method to subgroup analysis, and in particular to whether affect aggregates can be extracted with the same quality for different subpopulations, even if sampling error is overcome. To facilitate such efforts, we publish our Python code under MIT license [^4] https://github.com/dess-mannheim/temporal-adapters.
@@
YouGov pays its panel members adequately for their participation in surveys and is a member of several market research organisations (British Polling Council, ESOMAR, MRS) whose standards it adheres to (YouGov 2024b). We work with social media data that was publicly available at the time that we gathered it. We do not publish raw text data, or the Temporal Adapters that we trained, to protect the privacy of the individuals whose tweets are included in our training data. We do, however, publish the code that is needed to replicate our results under MIT license to support future research. Our method relies on more training data than the vast majority of social media users produce in a week, so it only works on an aggregate level. This limits the risks associated with profiling individuals through our method, even if we cannot fully rule out unethical applications. Our main results (Figure 4) highlight the importance of validation and comparison to survey data for each context and construct of interest. Still, an application of our method without such rigor could inform wrong or discriminatory downstream decisions, either because of bias in training data or because of measurement error.
Conclusion
This paper expands the inventory of methods for affect analysis available to our research community. Our work closes a temporal misalignment gap in previous surveys with LLMs by proposing Temporal Adapters that are trained on longitudinal social media data. To the best of our knowledge, our paper is the first to extend the analysis of affect in LLMs to a longitudinal setting. Temporal Adapters for LLMs open up new ways for studying affect aggregates in social media data longitudinally. More broadly, they may also enable future, more temporally-aligned LLM-based studies of human attitudes, values, and opinions.
Appendix A Appendix
@@
Figure 10: Time Series of Attitudes towards the NHS from Temporal Adapters. We extract answer probabilities by prompting a weekly fine-tuned Llama 3 8B with the same question wording as in the survey (YouGov 2024c), and compare them to the respective weekly survey data. The time series are min-max normalized and a $3$ week rolling average is applied. The shaded orange area indicates minimum and maximum LLM answer probabilities across 3 training seeds. YouGov survey data for comparison is only available in monthly waves.
@@
Figure 11: Results from Partial Hyperparameter Search. We synthetically mix LLM training data with splits ranging from data that is labeled $100\%$ sad to $100\%$ happy. We then extract answers to YouGov (2024a)’s survey question at each split, and show mean and standard deviation over 10 training seeds. Unless otherwise noted, each plot shows results after 50 training steps, with learning rate $5*10^{-6}$, temperature $1$, and using an answer prefix for answer extraction. We aim at a linear relationship between training mix and extracted answers, with low random error across training seeds.

### v2

@@ 5 Results
We implement two additional survey questions in addition to the one used in Britain’s Mood, Measured Weekly. First, we demonstrate the robustness of our extraction method and extract the collective emotions scared and sad using the PANAS-X survey instrument. The instrument works quite differently from YouGov’s questionnaire in that it measures each emotion with multiple adjectives, with the same answer options for each adjective: ‘very slightly or not at all’, ‘a little’, etc. Again, we cross-correlate the extracted collective emotions with the YouGov survey data – Figure 6 shows the results from 3 training seeds across the different extraction methods. We find that both scared and sad achieve cross-correlations comparable to the ones we extracted with the original YouGov question wording. This supports the robustness of our extraction method across prompt formulations, i.e., survey instruments, when measuring the same construct.
(a) Britain’s Mood: Scared
(b) Britain’s Mood: Sad
Figure 6: Robustness across multiple survey instruments. We extract the same collective emotions by querying Llama 3 with an additional survey instrument (PANAS-X: Clark and Watson 1994) and cross-correlate our results with the respective YouGov survey data. Across $3$ training seeds, we indicate minimum and maximum correlation with error bars and show the worst $p$ value (**$p<0.01$,***$p<0.001$). For time series, see Figure 11 in the Appendix.
(a) The NHS will…
(b) Boris Johnson is…
(c) Government Approval
Figure 7: Successful Application to Additional Questions of Public Opinion Beyond Collective Emotions. We extract collective attitudes towards the National Health Service (NHS), Boris Johnson as a Prime Minister, and the UK Government using survey questions developed by YouGov (2024e, c, b). We then cross-correlate our results with the respective YouGov survey data. Across $3$ training seeds, we indicate minimum and maximum correlation with error bars and show the worst $p$ value (*$p<0.05$,***$p<0.001$). For time series, see Figures 12, 13, and 14 in the Appendix. Our method can be used to extract collective answers to various issues of public opinion beyond collective emotions, as opposed to existing methods (such as TweetNLP), which have to be individually trained for specific questions.
@@
Second, we go beyond measuring collective emotions and test a question that measures affective stance towards the National Health Service (NHS), towards Boris Johnson as a Prime Minister, and towards the UK Government in general. YouGov runs most of its ‘trackers’ only in monthly waves, including on the attitude towards the NHS and towards Boris Johnson. This means that during the time period that we study, we only have 9 data points available. This shows the benefits of extracting attitudes from social media data, an approach that is orders of magnitude cheaper as well as of higher temporal resolution than traditional surveys. Figure 7 shows the cross-correlation of the answer options for the 9 available data points between our estimates and the survey data – for time series, see Figures 12, 13, and 14 in the Appendix. We find a strong positive cross-correlation between our estimate of government disapproval and the YouGov survey data. We also find moderate positive correlations for attitudes towards the NHS and towards Boris Johnson as a Prime Minister, given that we only have 9 data points available for evaluation. Existing NLP methods with pre-trained classifiers (such as TweetNLP) or dictionaries are always only suitable for extracting few, selected affect aggregates. This highlights a main benefit of our method compared to existing NLP methods, as it can be flexibly applied to answer any survey question without the need to re-train the Temporal Adapters.
6 Discussion
Contribution
This paper expands the inventory of methods for affect analysis available to our research community. We present a novel method for extracting longitudinal affect aggregates by using Temporal Adapters fine-tuned on user-generated data from social media. Our main focus lies on studying collective emotions in times of crisis, in particular, during the COVID-19 pandemic. Research on collective emotions has been shown to provide actionable insight for public health messaging and to study the social resilience of communities. We extend methods that have been previously used in this research area by proposing to extract survey answers with Temporal Adapters, rather than annotating and training specialized classification models. The main benefit of our approach is that is can be flexibly used with any survey question. This allows researchers to make use of the wealth of established survey questionnaires to extract affect aggregates from social media data, potentially from people who are inaccessible to traditional survey research. Ideally, such endeavors would use social media data from a clearly defined sample of users that represent the target population of a study. In the future, our method could allow for more timely estimates of affect aggregates and of public opinion, both within times of crisis and beyond that.
Our work closes a temporal misalignment gap in previous surveys with LLMs by proposing Temporal Adapters that are trained on longitudinal social media data. To the best of our knowledge, our paper is the first to extend the analysis of affect in LLMs to this longitudinal setting. Our approach creates novel ways to study affect aggregates in social media data longitudinally. More broadly, they may also enable future, more temporally-aligned LLM-based studies of human attitudes, values, and opinions.
@@ 6 Discussion
Our method focuses on the extraction of separate time series data, rather than comparing the prevalence of emotions or attitudes cross-sectionally. Our motivation behind that is that the relative probabilities of answer options in a cross-sectional sense solely depend on LLM pre-training. Future research should combine these, so far rather disparate, approaches by correcting for differences that stem from model pre-training. The results we present were obtained by training weekly Temporal Adapters, following exactly the intervals in which the respective survey waves were conducted. This enables comparability between our estimates and survey data, but even higher temporal resolution than usually used in surveys, a potential benefit of using social media data, should be investigated in future research. In contexts where survey data is available, survey data could in the future be used as a prior for predicting affect aggregates with higher temporal resolution from social media data, i.e. filling the gaps in between waves. At the same, our method also offers a lot of potential in contexts in which no (recent) survey data is available, for diverse reasons such as lack of infrastructure, an ongoing violent conflict or similar factors leading to extreme non-response. Like any other attitude extraction method, our approach requires validation in novel contexts, with possible design decisions ranging from model training hyperparameters and data sets used to the exact question wording and the answer extraction method. Application to additional contexts and points in time is particularly important since we base our analysis on data collected during the unique situation of the COVID-19 pandemic. Future research should further investigate the applicability of this method to subgroup analysis, and in particular to whether affect aggregates can be extracted with the same performance for different subpopulations, even if sampling error is overcome. To facilitate such efforts, we publish our Python code under MIT license. [^4] https://github.com/dess-mannheim/temporal-adapters
@@
YouGov pays its panel members adequately for their participation in surveys and is a member of several market research organisations (British Polling Council, ESOMAR, MRS) whose standards it adheres to (YouGov 2024d). We work with social media data that was publicly available at the time that we gathered it. We do not publish raw text data, or the Temporal Adapters that we trained, to protect the privacy of the individuals whose tweets are included in our training data. We do, however, publish the code that is needed to replicate our results under MIT license to support future research. Our method relies on more training data than the vast majority of social media users produce in a week, so it only works on an aggregate level. This limits the risks associated with profiling individuals through our method, even if we cannot fully rule out unethical applications. Our main results (Figure 4) highlight the importance of validation and comparison with survey data for each context and construct of interest. Still, an application of our method without such rigor could inform wrong or discriminatory downstream decisions, either because of bias in training data or because of measurement error.
Appendix
@@
Figure 10: Time Series of Emotions from Temporal Adapters, Part 3. We extract answer probabilities by prompting a weekly fine-tuned Llama 3 8B with the same question wording as in the survey (YouGov 2024a), and compare them to the respective weekly survey data. The time series are min-max normalized and a 3 week rolling average is applied. The shaded orange area indicates minimum and maximum LLM answer probabilities across 3 training seeds.
@@
Figure 12: Estimates of Attitudes towards the National Health Service from Temporal Adapters. We extract answer probabilities by prompting a weekly fine-tuned Llama 3 8B with the same question wording as in the survey (YouGov 2024e), and compare them to the respective weekly survey data. The time series are min-max normalized and a 3 week rolling average is applied. The shaded orange area indicates minimum and maximum LLM answer probabilities across 3 training seeds. YouGov survey data for comparison is only available in monthly waves. To the best of our knowledge, no pre-trained classifier is available for labeling tweets with regards to expressed attitudes towards the NHS.
Figure 13: Estimates of Attitudes towards Boris Johnson as Prime Minister from Temporal Adapters. We extract answer probabilities by prompting a weekly fine-tuned Llama 3 8B with the same question wording as in the survey (YouGov 2024c), and compare them to the respective weekly survey data. The time series are min-max normalized and a 3 week rolling average is applied. The shaded orange area indicates minimum and maximum LLM answer probabilities across 3 training seeds. YouGov survey data for comparison is only available in monthly waves. To the best of our knowledge, no pre-trained classifier is available for labeling tweets with regards to this survey question.
Figure 14: Estimates of Government Approval from Temporal Adapters. We extract answer probabilities by prompting a weekly fine-tuned Llama 3 8B with the same question wording as in the survey (YouGov 2024b), and compare them to the respective weekly survey data. The time series are min-max normalized and a 3 week rolling average is applied. The shaded orange area indicates minimum and maximum LLM answer probabilities across 3 training seeds. To the best of our knowledge, no pre-trained classifier is available for labeling tweets with regards to this survey question.
Figure 15: Britain‘s Mood: Pearson cross-correlations between the YouGov survey data and our estimates, by LLM temperature. The shaded orange area indicates minimum and maximum cross-correlation across 3 training seeds. We observe that the cross-correlation of many, but not all, answer options is stable across a wide range of LLM temperatures. In this paper, we estimate survey answers based on token probabilities and, unless otherwise specified, use the Llama 3 default temperature of $0.6$ for answer extraction.
@@
Figure 16: Results from the Partial Hyperparameter Search. We synthetically mix LLM training data with splits ranging from data that is labeled $100\%$ sad to $100\%$ happy. We then extract answers to YouGov (2024a)’s survey question at each split, and show mean and standard deviation over 10 training seeds. Unless otherwise noted, each plot shows results after 50 training steps, with learning rate $5*10^{-6}$, temperature $1$, and using an answer prefix for answer extraction. We aim at a linear relationship between training mix and extracted answers, with low random error across training seeds.